OpenStudy (anonymous):

Simplify the following expression: a + 2c - 5b - c + a - b

OpenStudy (anonymous):

Simplify or solve?

OpenStudy (anonymous):

simplify

OpenStudy (anonymous):

is it multiple choice?

OpenStudy (anonymous):

2a - 6b + 2c 2a - 6b + c a - 6b + c a - 5b + 2c

OpenStudy (anonymous):

2a - 6b + c

OpenStudy (anonymous):

thank yoy

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

you work for open study

OpenStudy (anonymous):

no just a user whose really good at math

OpenStudy (anonymous):

ahaha ok you can do another one?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Evaluate the following expression using the given values: Find x - 3y if x = 3 and y = -2. -9 -3 3 9

OpenStudy (anonymous):

Thats C, 3.

OpenStudy (anonymous):

THANK YOU!

OpenStudy (anonymous):

can you keep going

OpenStudy (anonymous):

Lol your welcome. Okay just 2 more

OpenStudy (anonymous):

Choose the best definition for the following term: substitution. A letter that holds the place for some unknown value in mathematics A process where you must look for terms that have identical variable parts and then combine their coefficients A process where if two things are equal, one can be put in the place of the other and nothing will change Terms that have identical variable parts

OpenStudy (anonymous):

C, A process where if two things are equal, one can be put in the place of the other and nothing will change,

OpenStudy (anonymous):

If the problem is 10 - 10 you can switch the 10s, into eachothers spots, and the answer will still be 0

OpenStudy (anonymous):

wow your really smart

OpenStudy (anonymous):

1 more

OpenStudy (anonymous):

Thanks. okay shoot

OpenStudy (anonymous):

Evaluate the following expression using the values given: Find 3x2 - y3 - y3 - z if x = 3, y = -2, and z = -5.

OpenStudy (anonymous):

I don't know that one sorry.

OpenStudy (anonymous):

its cool neither do i lol